OpenStudy (anonymous):

How do I simplify e^[-4 ln (tanϴ)]?

OpenStudy (anonymous):

Hi @baru. Any idea? :) I have a feeling that we should be focusing on the -4 ln (tanϴ) part but I'm not sure...

OpenStudy (baru):

consider z=ln( tanϴ) then e^z=tan(ϴ)

OpenStudy (anonymous):

I'm still a bit confused, I'm really not sure what to do!

OpenStudy (baru):

\[e^{-4\ln(\tan \theta)}=e^{\ln(\tan^{-4} \theta)}\]

OpenStudy (baru):

\[Let ~ \ln(\tan^{-4}\theta)=z\\ e^z=(\tan^{-4}\theta)\]

OpenStudy (anonymous):

Okay I understand it so far :)

OpenStudy (baru):

\[e^{\ln(\tan^{-4} \theta)}\] substitute 'z' in the above equation wherever possible

OpenStudy (baru):

so we just get \[e^z\]

OpenStudy (anonymous):

I'm just confused because can't we just do ln[tan(theta)]/e^4?

OpenStudy (baru):

no you cant\[a^{-bc}\neq \frac{c}{a^b}\] what i've typed above is what you tried to do

OpenStudy (anonymous):

Oh I see.

OpenStudy (anonymous):

Can I read my notes again and come back later? Thanks for all your help by the way. I just want to make sure I really understand this content!

OpenStudy (baru):

sorry if i confused you... the main thing to remember is this rule \[e^{\ln(a)}=a\]

OpenStudy (anonymous):

Thanks, I'll look into it. And don't worry, perhaps I just need to make sure I really understand this material! :)

OpenStudy (baru):

sure :)

OpenStudy (kenshin):

Baru I just typed both e^(-4lntan(x)) and (tan(x))^-4 into my graphing calculator; the graphs are slightly different, it appears whilst certain parts are the same, the exponent version has some parts missing, is this because there are domain restrictions for x inside ln?

OpenStudy (baru):

hmm curious... i entered the starting expression into wolframalpha, the answer seems correct

OpenStudy (kenshin):

tan(x) cannot be between (-infinity, 0) whilst inside the LN?

OpenStudy (kenshin):

Wolfram seems to have same graph for both for me too.

OpenStudy (baru):

yea ur right, tan cannot take negative values here, in any case, this is an exercise to get one comfortable with laws of exponents and logarithms.. so OP need not worry :)

OpenStudy (baru):

tan will take negative values in the 2nd and 4th quadrants, so i guess parts of the graph will go missing periodically

OpenStudy (baru):

then again, a negative number raised to an even power is positive... so perhaps you typed it slightly differently in wolfram and in your calculator

OpenStudy (baru):

the question exactly the way it appears is undefined in some parts however in \[e^{\ln(\tan^{-4} \theta)}\] tan is raised to an even power

OpenStudy (kenshin):

ok that's right, so tan(x) cannot be zero only, negative values can be valid. So x cannot be k(pi) where k = integers and 0, all other values are ok.

OpenStudy (baru):

i guess... but i'm no expert :p